User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!


Hobart 12m:

Disk VSN: HOB+0118

Data volume at beginning: 13354.874 GB

  • 1830 UT Experiment started OK (Ross)
  • 11:15 UT - PCFS computer ran out of room for the log file (again). Missing all log entries up until 11:32 UT. (Ellen)

Katherine 12m:

Disk VSN: NAIC-004

Data volume at beginning: 0.000 GB

  • 1830 UT Experiment started OK (Ross)

Yarragadee 12m:

Disk VSN: UVLBI-36

Data volume at beginning: 0.000 GB

  • 1830 UT Experiment started OK (Ross)
  • * 1908 UT power outage due to severe storms; dbbc had to be restarted twice on site. Google Chrome on ops2 was also helpfully unresponsive too for some reason, too little memory still?? Missed scans 098-1909 through 098-1943 (Ross)
  • 2110 UT another power outage, dbbc restarted again. Missed scans 098-2111 through 098-2115 (Ross)
  • 2221 UT dbbc restarted again. Missed scan 098-2223 (Ross)
  • 2326 UT GPS clock monitoring was not running on timing PC. Restarted now (Lucia)
  • 2326 UT data lagging 67GB behind at the moment (Lucia)
  • 0046 UT monitoring alarm because of break between scans of >15min (Lucia)
  • 0155 UT dbbc died again due to power glitches in Yarragadee in bad weather. Missed scans 099-0129 through 099-0148. First good scan (though a bit late on source) is 099-0152. If this happens again, we intend to change the battery - Randall is looking for one at the moment. (Lucia)
  • 0250 UT dbbc failure due to power irregularities. New backup power module (UPS) is in now. Missed scans: 099-0232, 099-0235, 099-0238. (Lucia)
  • 0702 UT realised that the GPS receiver is not working properly. Antenna controller clock changed to the Hobart GPS clock. Run fmset several times. Then ACU time difference was huge (up to 8 sec). Fixed itself once the antenna had the new clock. In fmset, clocks are synchronised but 0.1 sec offset (how to fix (that?). Also the FS was restarted. Affected scans: ~ 06:30UT until 099-0707. First good scan 099-0712. (Lucia)
  • 09:16 UT - Occasionally checking the fmset, but there is still a consistent offset between the mark 5 and the other 2 of either 0.1 or 0.2 sec. Combined efforts of Jim, Jamie and Lucia could not fix this, so are continuing in the hope that it might fix itself. (Ellen)
  • 09:16 UT - Disc pos ~ 138 GB behind due to missed scans. (Ellen)
  • 09:33 UT - Tsys bands 09, 10, 12 slightly too high, bands 13, 14 way too high, though storm seems to have passed. (Ellen)
  • 13:13 UT - Checked the clocks and they are now synced, not sure exactly when this occurred - somewhere in the last 2 hours. (Ellen)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r4734.txt · Last modified: 2016/04/08 18:58 by Lucia McCallum